# Non-thermal X-ray Emission from Merging Massive Black Hole Binaries

### v1

@@ Abstract
Recent hydrodynamical simulations have identified a disappearing thermal X-ray signature in massive black hole binaries (MBHBs) embedded in circumbinary disks, arising from the tidal truncation and depletion of minidiscs shortly before merger. This feature has been proposed as a promising electromagnetic counterpart to MBHB mergers detectable by LISA. In this work, we examine whether non-thermal X-ray emission powered by magnetic reconnection could obscure or modify this thermal X-ray drop. We construct semi-analytic models for both the thermal X-ray emission from minidiscs and the non-thermal synchrotron emission produced by reconnection in magnetically dominated black hole magnetospheres. Evaluating these models across the MBHB mass range relevant for LISA, we find that for physically motivated magnetic field strengths and accretion rates, the non-thermal X-ray luminosity remains several orders of magnitude below the thermal component throughout the inspiral. Even under optimistic assumptions that enhance the non-thermal emission, it remains significantly subdominant. We further incorporate the magnetospheric balding framework to model the decay of non-thermal emission near merger, finding that reconnection-powered X-ray emission fades on short, mass-scaled timescales once the external magnetic flux supply is disrupted. Taken together, our results indicate that non-thermal emission is unlikely to mask the disappearing thermal X-ray signature, reinforcing its robustness as an electromagnetic counterpart to MBHB mergers and its potential utility for multi-messenger studies with LISA.
@@ 1 Introduction
During the very late inspiral and early post-merger of a MBHB+CBD system, the progressive depletion of the minidiscs and accompanying decline in accretion rate may naturally drive the system toward such a magnetically dominated state (Most & Wang, 2024). Global GRMHD simulations of MADs show that strongly magnetized flows generically develop thin current sheets near the black holes, where magnetic reconnection can efficiently convert magnetic energy into non-thermal particle acceleration and high-energy radiation (Ripperda et al., 2020; Chashkina et al., 2021; Ripperda et al., 2022). This process is capable of producing luminous synchrotron emission extending into the X-ray and soft-$\gamma$-ray bands.
As the system evolves through merger and the inflow of magnetic flux diminishes, reconnection can rapidly dissipate the large-scale magnetic fields threading the black holes, leading to a “balding” of their magnetospheres (Bransgrove et al., 2021). The associated high-energy emission would then decay on short timescales, potentially producing a brief non-thermal flare followed by rapid fading.
@@
Our main result suggests that, across all realistically feasible regions of parameter space, the non-thermal X-ray component from synchrotron radiation should be several orders of magnitude lower than the emission from the thermal X-ray signature; hence, the disappearing thermal X-ray signature should still survive as a viable means for identifying host galaxies of MBHB mergers.
@@ 2.1 Thermal X-ray Emission from Minidiscs
We compute the total thermal X-ray luminosity by constructing a smooth, semi-analytic model spectrum whose shape and normalization are informed by the hydrodynamical simulations of Krauth et al. (2023). The simulated spectra motivate both the characteristic thermal peak and the overall spectral shape of the emission originating from the minidiscs. We then scale this spectrum with the total binary mass, allowing the peak energy and amplitude to shift according to physically motivated mass dependences consistent with thin-disk emission, while optionally permitting spin-dependent corrections to the radiative efficiency (though spin is not included in our fiducial results).
@@ 2.2 Non-thermal synchrotron model from magnetic reconnection
of the Blandford–Znajek jet power $L_{\rm BZ}$ is dissipated through magnetic reconnection and radiated, with $f_{\rm rec}$ taken to be an order-unity efficiency consistent with kinetic reconnection studies and the finding that radiative reconnection converts an $\mathcal{O}(10\%)$ fraction of the Poynting flux into radiation (Sironi & Spitkovsky, 2014; Hakobyan et al., 2023). The corresponding non-thermal X-ray luminosity is obtained by integrating the spectrum over the same $E_{\rm X\text{-}ray}=0.124$$124.0\penalty 10000\ \mathrm{keV}$ band. We evaluate the reconnection-powered spectrum across the MBHB mass range by computing $B(M,\dot{M})$ and the corresponding $L_{\rm BZ}$ at each point in a $(M,\dot{M})$ grid.
@@
In magnetically arrested accretion flows, the near-horizon magnetic field strength $B$ can be related to the dimensionless horizon magnetic flux $\phi_{B}$; however, we retain $B$ as an explicit parameter here to remain consistent with the calibrated reconnection–synchrotron framework adopted from kinetic simulations. Magnetic field strengths and accretion rates are drawn from physically motivated ranges expected for MAD-like accretion flows, and variations around a fiducial model are used to bracket plausible extremes. The explicit equations and fiducial parameter choices entering this procedure are summarized in Appendix § A.
@@ 2.3 Magnetospheric balding and decay of non-thermal emission
To model the decay of the non-thermal X-ray emission once the external magnetic flux supply diminishes, we adopt the magnetospheric “balding” framework developed by Bransgrove et al. (2021). In this picture, magnetic flux threading the black hole event horizon is removed through plasmoid-mediated magnetic reconnection once the external supply of flux diminishes. General-relativistic kinetic and resistive MHD simulations show that the horizon-threading magnetic flux decays quasi-exponentially in time, with a characteristic e-folding timescale set by the reconnection rate in the equatorial current sheet (Bransgrove et al., 2021; Ripperda et al., 2020; Ripperda et al., 2022).
@@ 3 Results
Figure 1 compares the mass scaling of the thermal X-ray luminosity from minidiscs to the expected non-thermal synchrotron contribution powered by magnetic reconnection. For fiducial, physically motivated choices of the magnetic field strength and accretion rate, the non-thermal synchrotron component lies several orders of magnitude below the thermal X-ray emission across the full MBHB mass range relevant for LISA. Even when the magnetic field strength and accretion rate are increased by an order of magnitude relative to the fiducial model—values that likely represent optimistic, if not extreme, conditions for magnetospheric reconnection—the resulting synchrotron luminosity remains suppressed by $\gtrsim 2$ orders of magnitude compared to the thermal component. Such a separation suggests that the disappearance of the thermal X-ray signal should remain observable and not be obscured by non-thermal emission, although a residual synchrotron contribution at these enhanced levels could in principle be detectable with sensitive X-ray monitoring.
Figure 1: X-ray luminosity $L_{\mathrm{X}}$ as a function of black hole mass for thermal (blue) and synchrotron (green) emission models. Solid and dashed blue curves show the thermal scaling for non-spinning ($a=0$) and rapidly spinning ($a=0.99$) black holes, respectively. The green solid curve denotes the fiducial synchrotron model, while dotted and dashed curves illustrate variations with magnetic field strength and accretion rate scaled by factors of 1010 and $10^{-1}$ relative to the fiducial values.
We note that our fiducial thermal model neglects black hole spin, while the non-thermal synchrotron emission is expected to be most efficient for rapidly spinning black holes that can sustain strong magnetospheres. Including spin-dependent corrections to the thermal radiative efficiency, as illustrated by the $a=0.99$ thermal model in Fig. 1, further increases the thermal X-ray luminosity and thus accentuates the contrast between the thermal and non-thermal components. This strengthens the conclusion that, even in systems hosting rapidly spinning black holes, the reconnection-powered synchrotron emission is unlikely to mask the pronounced thermal X-ray drop predicted to occur in the final stages of MBHB inspiral.
@@
Figure 2 shows the characteristic timescale on which non-thermal X-ray emission powered by magnetic reconnection is expected to decay once the supply of magnetic flux to the black hole magnetosphere is disrupted. Following the magnetospheric balding framework of Bransgrove et al. (2021), the luminosity fades quasi-exponentially with an e-folding time $\tau\sim 100$$500\,r_{g}/c$. Because the gravitational timescale scales linearly with black hole mass, the corresponding decay time increases proportionally across the MBHB mass range, from minutes for $M\sim 10^{4-5}M_{\odot}$ to days or longer for the most massive systems considered here ($M\sim 10^{9-10}M_{\odot}$). Also shown in Fig. 2 is an illustrative timescale for the recovery of thermal X-ray emission following merger. While the precise moment at which the thermal component can be considered “recovered” is not uniquely defined, this curve indicates the time required for the thermal X-ray luminosity to rebound to within approximately two orders of magnitude of its pre-merger level, comparable to even the most optimistic non-thermal contribution considered here. This recovery time remains significantly longer than the reconnection-driven decay time across the full black hole mass range.
@@
Figure 2: Time required for the non-thermal X-ray luminosity to decay by one e-fold due to magnetospheric balding, shown as a function of black hole mass. Coloured curves correspond to decay timescales $\tau=100\,r_{g}/c$, $300\,r_{g}/c$, and $500\,r_{g}/c$. The black dash–dotted curve indicates the time required for the thermal X-ray luminosity to recover to within two orders of magnitude of its pre-merger level.
As shown in Fig. 1, even pushing the upper bound of what is physically reasonable, the reconnection-powered synchrotron luminosity remains well below the thermal X-ray emission from the minidiscs. It can however still be instructive to ask how extreme the physical conditions must be for the reconnection-powered synchrotron emission to approach or exceed the thermal X-ray luminosity from the minidiscs. In our framework, this would require both magnetospheric field strengths to be tens of times what is expected for magnetically arrested accretion flows and accretion rates a few to ten times the Eddington limit. Such conditions correspond to an optimistic and likely short-lived corner of parameter space, particularly for MBHBs in the late inspiral phase where the minidiscs are simultaneously being depleted. Even under these assumptions, as Fig. 2 indicates, the non-thermal emission would decay rapidly once the external magnetic flux supply diminishes. Consequently, any reconnection-powered non-thermal emission that might transiently compete with the thermal signal before merger is likewise expected to fade, and on a comparatively short timescale, leaving the system X-ray faint in both channels until accretion and magnetic flux are re-established.

### v2

@@ Abstract
Recent hydrodynamical simulations have identified a disappearing thermal X-ray signature in massive black hole binaries (MBHBs) embedded in circumbinary disks, arising from the tidal truncation and depletion of minidiscs shortly before merger. This feature has been proposed as a promising electromagnetic counterpart to MBHB mergers detectable by LISA. In this work, we examine whether non-thermal X-ray emission powered by magnetic reconnection could obscure or modify this thermal X-ray drop. We construct semi-analytic models for both the thermal X-ray emission from minidiscs and the non-thermal synchrotron emission produced by reconnection in magnetically dominated black hole magnetospheres. Evaluating these models across the MBHB mass range relevant for LISA, we find that for physically motivated magnetic field strengths and accretion rates, the non-thermal X-ray luminosity remains several orders of magnitude below the thermal component throughout the inspiral, particularly in the soft X-ray band where the thermal emission is concentrated. Even under optimistic assumptions that enhance the non-thermal emission, it remains insufficient to erase the characteristic thermal drop, though a transient hard X-ray enhancement may arise near merger. We further incorporate the magnetospheric balding framework to model the decay of non-thermal emission near merger, finding that reconnection-powered X-ray emission fades on short, mass-scaled timescales once the external magnetic flux supply is disrupted. Taken together, our results indicate that non-thermal emission is unlikely to mask the disappearing thermal X-ray signature, reinforcing its robustness as an electromagnetic counterpart to MBHB mergers and its potential utility for multi-messenger studies with LISA.
@@ 1 Introduction
During the very late inspiral and early post-merger of a MBHB+CBD system, the progressive depletion of the minidiscs and accompanying decline in accretion rate may naturally drive the system toward such a magnetically dominated state (Most & Wang, 2024; Manikantan & Paschalidis, 2025). Global GRMHD simulations of MADs show that strongly magnetized flows generically develop thin current sheets near the black holes, where magnetic reconnection can efficiently convert magnetic energy into non-thermal particle acceleration and high-energy radiation (Ripperda et al., 2020; Chashkina et al., 2021; Ripperda et al., 2022). This process is capable of producing luminous synchrotron emission extending into the X-ray and soft-$\gamma$-ray bands.
As the system evolves through merger and the inflow of magnetic flux diminishes, reconnection can rapidly dissipate the large-scale magnetic fields threading the black holes, leading to a “balding” of their magnetospheres (Lyutikov, 2011; Lyutikov & McKinney, 2011; Bransgrove et al., 2021). The associated high-energy emission would then decay on short timescales, potentially producing a brief non-thermal flare followed by rapid fading.
@@
Our main result suggests that, across all realistically feasible regions of parameter space, the non-thermal X-ray component from synchrotron radiation remains several orders of magnitude below the thermal emission throughout the inspiral. Even when time-dependent spectral evolution and band-specific hard X-ray enhancement is taken into account, the non-thermal contribution strengthens primarily at higher photon energies and only after the thermal luminosity has already declined substantially. Consequently, the disappearing thermal X-ray signature persists as a robust and distinguishable feature, reinforcing its viability as a means for identifying host galaxies of MBHB mergers.
@@ 2.1 Thermal X-ray Emission from Minidiscs
We compute the total thermal X-ray luminosity by constructing a smooth, semi-analytic model spectrum whose shape and normalization are informed by the hydrodynamical simulations of Krauth et al. (2023). The simulated spectra motivate both the characteristic thermal peak and the overall spectral shape of the emission originating from the minidiscs. We then scale this spectrum with the total binary mass, allowing the peak energy and amplitude to shift according to physically motivated mass dependences consistent with thin-disk emission, while optionally permitting modest spin-dependent corrections to both the radiative efficiency and the effective inner disc radius.
@@ 2.2 Non-thermal synchrotron model from magnetic reconnection
of the Blandford–Znajek jet power $L_{\rm BZ}$ is dissipated through magnetic reconnection and radiated, with $f_{\rm rec}$ taken to be an order-unity efficiency consistent with kinetic reconnection studies and the finding that radiative reconnection converts an $\mathcal{O}(10\%)$ fraction of the Poynting flux into radiation (Sironi & Spitkovsky, 2014; Hakobyan et al., 2023). The corresponding non-thermal X-ray luminosity is obtained by integrating the spectrum over a specified energy range. In addition to the full X-ray band considered for our baseline comparison, we evaluate the synchrotron contribution across instrument-motivated soft and hard X-ray sub-bands to assess its observational detectability. We evaluate the reconnection-powered spectrum across the MBHB mass range by computing $B(M,\dot{M})$ and the corresponding $L_{\rm BZ}$ at each point in a $(M,\dot{M})$ grid.
Because the synchrotron spectrum is constructed explicitly as a function of magnetic field strength and accretion rate, this framework allows us to compute not only band-integrated luminosities but also time-dependent spectra and band-limited light curves as the binary separation decreases toward merger. In practice, we implement this through a smooth, separation-dependent scaling of the available reconnection power, motivated by studies that find increasing magnetic stress and dissipation as compact binaries approach coalescence (e.g., Hakobyan et al. 2023; Manikantan & Paschalidis 2025) and by the rapid decline in minidisc accretion rates seen in our simulations during the final stages of inspiral. We emphasize that this prescription is intended to bracket physically plausible behaviour rather than to provide a first-principles prediction of the exact merger light curve or detailed spectral evolution; our goal is to test whether any reasonable enhancement of reconnection-powered emission could plausibly obscure the thermal X-ray disappearance.
@@
In magnetically arrested accretion flows, the near-horizon magnetic field strength $B$ can be related to the dimensionless horizon magnetic flux $\phi_{B}$; however, we retain $B$ as an explicit parameter here to remain consistent with the calibrated reconnection–synchrotron framework adopted from kinetic simulations. Magnetic field strengths and accretion rates are drawn from physically motivated ranges expected for MAD-like accretion flows, and variations around a fiducial model are used to bracket plausible extremes. The explicit equations and fiducial parameter choices entering this procedure are summarized in Appendix A.
@@ 2.3 Magnetospheric balding and decay of non-thermal emission
To model the decay of the non-thermal X-ray emission once the external magnetic flux supply diminishes, we adopt the magnetospheric “balding” framework (Lyutikov, 2011; Lyutikov & McKinney, 2011), implementing the decay timescales following Bransgrove et al. (2021). In this picture, magnetic flux threading the black hole event horizon is removed through plasmoid-mediated magnetic reconnection once the external supply of flux diminishes. General-relativistic kinetic and resistive MHD simulations show that the horizon-threading magnetic flux decays quasi-exponentially in time, with a characteristic e-folding timescale set by the reconnection rate in the equatorial current sheet (Bransgrove et al., 2021; Ripperda et al., 2020; Ripperda et al., 2022).
@@ 3 Results
Figure 1 compares the mass scaling of the thermal X-ray luminosity from minidiscs to the non-thermal synchrotron contribution powered by magnetic reconnection. The upper shaded band shows the thermal emission integrated over the full X-ray range (0.124–124 keV), spanning the variation produced by black hole spin between $a=0$ and $a=0.99$. The two lower shaded bands show the reconnection-powered synchrotron luminosity, evaluated both over the full X-ray band (“Synch total”) and restricted to the soft X-ray range (0.124–12 keV). In each synchrotron case, the solid line indicates the fiducial model, while the surrounding shaded region brackets the “low” and “high” magnetic-field and accretion configurations described in Appendix A. Although the thermal spectrum is integrated over the full X-ray band for consistency, the emission is effectively confined below $\sim 12$ keV across the mass range considered. This makes the soft X-ray comparison particularly informative, since the full-band thermal luminosity is nearly identical to its soft-band contribution and most current and planned X-ray observatories operate over limited energy sub-bands rather than the entire X-ray range. At the highest black hole masses, the thermal band exhibits a downturn; this does not indicate an intrinsic suppression of minidisc emission, but instead reflects the shift of the thermal spectral peak to lower energies as $M$ increases, causing an increasing fraction of the emission to fall below the adopted X-ray bandpass.
For fiducial, physically motivated magnetic field strengths and accretion rates, the non-thermal synchrotron component remains well below the thermal X-ray emission across the LISA-relevant mass range. In the soft X-ray band, the synchrotron contribution remains subdominant even in the most optimistic “high” configuration ($B_{0}=10^{6}$ G, $\dot{M}=10^{-1}\dot{M}_{\rm Edd}$). When integrated over the full X-ray range, the total synchrotron luminosity can approach or exceed the thermal emission only at the highest masses and only when compared to the non-spinning ($a=0$) lower edge of the thermal band. However, the non-thermal synchrotron emission in our framework is most efficient for rapidly spinning black holes capable of sustaining strong magnetospheres, as the available Blandford–Znajek power increases with spin. Including spin-dependent corrections to the thermal radiative efficiency, as illustrated by the $a=0.99$ upper edge of the thermal band in Fig. 1, raises the thermal luminosity and restores a clear separation between the thermal and non-thermal components across the full mass range, regardless of whether the synchrotron emission is integrated over the soft or full X-ray band. Thus, even under extreme magnetic field strengths and accretion rates, the reconnection-powered synchrotron emission is unlikely to mask the pronounced thermal X-ray drop expected in the final stages of MBHB inspiral. We also note that the thermal luminosities shown here correspond effectively to face-on viewing geometries, consistent with the spectra presented in Krauth et al. (2023). Relativistic Doppler boosting at higher inclinations can enhance the high-energy thermal emission (see Appendix B of Krauth et al. (2023)), which would further increase the separation between the thermal and non-thermal components.
Figure 1: X-ray luminosity $L_{X}$ as a function of black hole mass. The upper shaded band (green) shows the range of thermal minidisc emission obtained by varying the black hole spin between $a=0$ and $a=0.99$, integrated over the full X-ray band (0.124–124 keV). The middle shaded band corresponds to the synchrotron luminosity powered by magnetic reconnection integrated over the full X-ray band (“Synch total”), while the lowest shaded band shows the synchrotron contribution restricted to the soft X-ray band (0.124–12 keV). In each synchrotron case, the solid curve denotes the fiducial model, and the surrounding shaded region spans the range between the “low” and “high” magnetic-field and accretion configurations. These correspond to $B_{0}=10^{4},\,10^{5},\,10^{6}\,\mathrm{G}$ and $\dot{M}=10^{-3},\,10^{-2},\,10^{-1}\dot{M}_{\mathrm{Edd}}$, respectively, using the magnetic-field scaling of Eq. (5). Overall, the thermal minidisc emission remains dominant across the LISA-relevant mass range for physically motivated parameters.
In addition to comparing band-integrated luminosities, we examine the time-dependent spectral evolution of both the thermal and non-thermal components, allowing us to assess how their relative contributions evolve across soft and hard X-ray energies as merger approaches. Figure 2 shows representative spectra evaluated at 100 hours before merger, 5 hours before merger, and at merger. For each epoch, we plot the thermal spectrum for both non-spinning ($a=0$) and rapidly spinning ($a=0.99$) black holes, together with the non-thermal synchrotron spectra for the fiducial and “high” magnetic field models. Vertical grey lines indicate the full X-ray range, broken into soft and hard bands.
At early times (100 hours before merger), the thermal component dominates the spectrum, notably in the soft X-ray band. The maximally spinning thermal model exhibits a higher peak luminosity and modestly harder spectrum relative to the non-spinning case, reflecting the increased radiative efficiency and smaller effective inner disc radius. In contrast, the synchrotron emission remains comparatively suppressed at large binary separations. This behaviour is consistent with reconnection-powered models in which the non-thermal luminosity strengthens as the binary separation decreases and magnetic stresses intensify (e.g., Manikantan & Paschalidis 2025). Even in the “high” magnetic field configuration, the non-thermal component contributes primarily at higher photon energies at these early stages.
By 5 hours before merger, the thermal emission has declined substantially as the minidiscs become increasingly truncated and depleted. The spectral peak shifts downward in luminosity, and the soft X-ray band begins to dim appreciably. Over the same interval, the synchrotron spectrum strengthens and extends to higher energies, with the “high” model becoming comparable to the thermal emission at hard X-ray energies while remaining subdominant in the soft band.
At merger, the thermal component is strongly suppressed across the X-ray range as the minidiscs effectively vanish. In this phase, the non-thermal synchrotron spectrum dominates the high-energy X-ray band, particularly for the “high” magnetic field configuration, while the fiducial model provides a more moderate but still significant high-energy contribution.
For the maximally spinning ($a=0.99$) thermal model, spin-dependent thin-disc corrections are applied only during the earlier, radiatively efficient phase. In our framework, we model the late inspiral and merger stages as transitioning toward a magnetically arrested disk (MAD) state, in which a significant fraction of the accretion power is redirected into large-scale magnetic fields and jet production rather than radiated thermally. In this regime, the assumptions underlying thin-disc radiative efficiency are no longer strictly applicable. We therefore suppress spin-dependent thermal enhancements at late times to avoid artificially boosting the thermal luminosity in a magnetically dominated phase where non-thermal processes are expected to play an increasingly important energetic role.
Figure 2: Spectral evolution of thermal and non-thermal emission during the final stages of MBHB inspiral. Shown are spectra at 100 hours before merger (left), 5 hours before merger (middle), and at merger (right). The peaked spectra (green) correspond to the thermal minidisc emission, with the solid curve showing the non-spinning case ($a=0$) and the dashed curve the maximally spinning case ($a=0.99$). The steadily rising power-law spectra (blue) show the non-thermal synchrotron emission powered by magnetic reconnection, where the solid line denotes the fiducial magnetic-field configuration and the dashed line represents the “high” magnetic-field model described in Appendix A. Vertical dotted lines indicate representative soft and hard X-ray energy bands. At early times the spectrum is dominated by thermal emission, particularly in the soft X-ray band, while toward merger the thermal component diminishes and the synchrotron contribution becomes increasingly important at higher energies.
To connect the evolving spectra to observational diagnostics more directly, we next compute the time-dependent luminosity in representative instrument-specific X-ray bands. This allows us to assess how the relative thermal and non-thermal contributions evolve within the energy ranges probed by current and future facilities.
Figure 3 shows the evolution of the band-integrated luminosity over the final 100 hours before merger across three representative X-ray energy bands. The left panel shows the soft X-ray band (0.10.11212 keV), relevant for current and forthcoming soft X-ray facilities such as Chandra, XMM-Newton, XRISM, eROSITA, and Athena mission (Weisskopf et al., 2002; Jansen et al., 2001; Tashiro et al., 2021; Predehl et al., 2021; Nandra et al., 2013). In this band the thermal emission dominates early and drops sharply as the minidiscs are depleted during the late inspiral. For fiducial parameters the synchrotron contribution remains several orders of magnitude lower throughout the entire inspiral, becoming comparable only near merger in optimistic configurations.
The middle and right panels show the harder X-ray bands probed by NuSTAR (337979 keV) and Swift-BAT (1515150150 keV), respectively (Harrison et al., 2013; Barthelmy et al., 2005). Because the thermal spectrum is effectively confined to photon energies below $\sim 12$ keV, its contribution to these higher-energy bands is negligible, leaving the emission largely governed by the non-thermal synchrotron component. Toward merger, this non-thermal luminosity gradually increases. While the strength of the hard-band enhancement depends sensitively on magnetic-field strength and accretion rate, its emergence near merger reflects the increasing dynamical importance of magnetic stresses as the system transitions toward a magnetically dominated state.
Taken together, these band-resolved light curves indicate that the final stages of MBHB inspiral may exhibit not only a luminosity decline, but a spectral reconfiguration, with soft X-rays tracing minidisc depletion and hard X-rays probing the growing influence of magnetospheric reconnection. While this enhancement is most pronounced in the hard X-ray bands and its observational detectability depends sensitively on source mass, distance, and exposure time, the resulting spectral transition may nevertheless provide an observational handle on the approach to merger, linking the soft X-ray disappearance with a potential hard-band signature of magnetospheric activity.
Figure 3: Time-dependent X-ray luminosity during the final $\sim 100$ hours before merger in representative instrumental bands. The left panel shows the soft X-ray band (0.1–12 keV), while the middle and right panels show the hard X-ray bands 3–79 keV and 15–150 keV, respectively. In the soft band, the solid curve denotes the thermal luminosity extracted from the hydrodynamical simulations. The non-thermal synchrotron contribution is shown in each panel by the lower curves: the solid line indicates the fiducial reconnection-powered model, while the surrounding shaded region brackets the range spanned by the “low” and “high” magnetic-field configurations described in Appendix A. Although the thermal emission dominates the soft X-ray band, it contributes negligibly at higher photon energies, leaving the hard X-ray bands largely governed by the non-thermal synchrotron component. As merger approaches, the thermal luminosity drops sharply while the synchrotron contribution gradually strengthens.
@@
The duration of this high-energy phase is constrained by the subsequent evolution of the magnetosphere. Figure 4 shows the characteristic timescale on which reconnection-powered non-thermal X-ray emission is expected to decay once the external magnetic flux supply to the black hole magnetosphere is disrupted. Following the magnetospheric balding framework of Bransgrove et al. (2021), the horizon-threading magnetic flux dissipates quasi-exponentially with an e-folding time $\tau\sim 100$$500\,r_{g}/c$. Because this gravitational timescale scales linearly with black hole mass, the corresponding decay time increases proportionally across the MBHB mass range, from minutes for $M\sim 10^{4-5}\,M_{\odot}$ to days for $M\sim 10^{9-10}\,M_{\odot}$. This decay sets the duration of any merger-associated hard X-ray enhancement.
Figure 4 also shows an illustrative recovery timescale for the thermal X-ray emission following merger. Although the precise definition of “recovery” is not unique, this curve marks the time required for the thermal luminosity to return to within two orders of magnitude of its pre-merger level—comparable to even the most optimistic non-thermal contribution considered here. Across the full mass range, this thermal recovery time exceeds the reconnection-driven decay time. Consequently, even in optimistic magnetic-field configurations where the non-thermal component briefly approaches the declining thermal luminosity at high photon energies, the reconnection-powered emission fades before the thermal emission rebounds. The system is therefore expected to enter a transient X-ray–dark phase following merger, until accretion and magnetic flux are re-established.
@@
Figure 4: Time required for the non-thermal X-ray luminosity to decay by one e-fold due to magnetospheric balding, shown as a function of black hole mass. The three lower curves correspond, from bottom to top, to decay timescales $\tau=100\,r_{g}/c$ (dotted), $300\,r_{g}/c$ (solid), and $500\,r_{g}/c$ (dashed). The upper dash–dotted curve shows an illustrative recovery timescale for the thermal X-ray emission, defined as the time required for the thermal luminosity to return to within two orders of magnitude of its pre-merger level. Across the full MBHB mass range, the non-thermal decay timescale remains shorter than the corresponding thermal recovery timescale, implying that any transient non-thermal enhancement fades before the thermal emission is re-established.
As shown in Fig. 1, only under the most extreme magnetic field strengths and accretion rates does the reconnection-powered synchrotron emission begin to compete with the thermal X-ray luminosity, and then primarily in the full X-ray band at the highest black hole masses. Reaching such levels in our framework requires magnetospheric field strengths significantly above those typically expected for magnetically arrested accretion flows and accretion rates approaching or exceeding the Eddington limit—conditions that are both optimistic and likely short-lived during the late inspiral phase when the minidiscs are simultaneously being depleted. Even in this regime, the balding-driven decay timescales shown in Fig. 4 imply that any transient non-thermal enhancement would fade rapidly once the magnetic flux supply is disrupted. Thus, while extreme configurations may briefly elevate the hard X-ray output, they do not prevent the system from entering the merger-associated X-ray–dark phase described above.